# Special Weingarten surfaces with planar convex boundary

### v1

@@ 5 Two Ros-Rosenberg type theorems for SW-surfaces
In the former case, we make a natural assumption on the function $g$ in order to get Ros-Rosenberg result for every surfaces obtained from surfaces in $\mathcal{W}_{g}$ after suitable homotheties, whose ratios can be computed explicitly. We point out that condition (5.1) implies that, when we do a homothety of a surface in $\mathcal{W}_{g},$ its curvature diagram does not intersect a line with slope $m_{0}$ passing through the origin. This allows us to compare our surface with a suitable catenoid by the maximum principle (see Section 1.3).
@@
In view of Remark 5.3, we may suppose that $\Gamma$ is not a circle, that is, $\Lambda>\lambda>0$. By convexity, any circle of radius $1/\lambda$ (resp. $1/\Lambda$) tangent to $\Gamma$ with inward unit normal vector pointing inside $\Omega$ must bound a closed domain containing $\Gamma$ (resp. contained in the closed domain bounded by $\Gamma$). We also define $\mathcal{C}$ as the smallest circle in $P$ enclosing $\Gamma$, whose center and radius we denote by $c_{\Gamma}$ and $\omega(\Gamma)$, respectively. In particular, $1/\Lambda\leq\omega(\Gamma)\leq 1/\lambda$.
@@
Let us fix a generic unit horizontal vector $v$ and let $Z$ be the horizontal cylinder in $\mathcal{W}_{f}$ whose revolution axis is the straight line orthogonal to $v$, contained in the plane $\{z=d/\beta\}$ and passing through $(c_{\Gamma},d/\beta)$. In this way, the plane $P$ is tangent to $Z$, and the maximum height of $Z$ with respect to $P$ is equals to $2d/\beta$. If $Q_{v}$ is a vertical plane orthogonal to $v$ passing through $(c_{\Gamma},d/\beta)$, we define $\widetilde{Z}$ as the part of $Z$ lying below $Q_{v}$ with respect to $v$, and $\widetilde{Z}(t):=\widetilde{Z}-tv$ as the image of $\widetilde{Z}$ by the translation with respect to the vector $-tv$.
For values $t\gg 0$, $\widetilde{Z}(t)$ does not touch $\Sigma$. By decreasing $t$, $\widetilde{Z}(t)$ approaches $\Sigma$ and by the maximum principle, there cannot be a first contact point between $\widetilde{Z}(t)$ and $\Sigma$ for $t>\omega(\Gamma)$. This implies that $\Sigma$ is contained in the upper (with respect to $v$) closed half-space bounded by $Q_{v}-(\omega(\Gamma)+d/\beta)v$. As $v$ was generically chosen, we deduce that $\Sigma$ is contained in $D((c_{\Gamma},0),\omega(\Gamma)+d/\beta)\times[0,h_{\Sigma}]$.
@@
This yields that $\omega(\Gamma)+d/\beta\leq 1/\lambda+d/\beta<d/\alpha$ and then, the cylinder $\partial S^{+}\times[0,+\infty)$ strictly contains the truncated solid cylinder $D((c_{\Gamma},0),\omega(\Gamma)+d/\beta)\times[0,h_{\Sigma}]$ in its interior. Define $S(\tau):=S^{+}+\tau e_{3}$ to be the translation of $S^{+}$ by the vector $\tau e_{3}$. For any $\tau>2d/\beta,$ the surfaces $S(\tau)$ and $\Sigma$ do not meet. By decreasing $\tau$, if there is a first contact point between $S(\tau_{0})$ and $\Sigma$, for some $\tau_{0}\geq 0$, such a contact point must be necessarily an interior tangency point, since the domain bounded by the orthogonal projection of the boundary of $S(\tau_{0})$ on $P$ strictly contains $\Gamma$. The existence of such an interior contact point is a contradiction by the maximum principle. Therefore, $\Sigma$ must be strictly below $S(0)$.
Since $1/\lambda<d/\alpha$, we can translate $S(0)$ horizontally in such a way that its boundary touch every point of $\Gamma$ we desire from outside. Again by the maximum principle, this implies that $\Sigma$ is contained in the solid cylinder $\overline{\Omega}\times[0,\infty)$. Now, it is easy to prove, using the Alexandrov reflection method with horizontal planes, that $\Sigma$ is a graph over $\Omega$ and therefore a topological disk.
@@
Let $p_{\max}\in\Sigma$ be a point at maximal height and let $\mathcal{O}$ denote the open region bounded by $\Sigma\cup\Omega$. By doing the Alexandrov reflection method with horizontal planes, we get that the reflection of $p_{\max}$ with respect any plane $P(t)$ for $h_{\Sigma}/2<t<h_{\Sigma}$ lies in $\mathcal{O}$. This means the reflection of $p_{\max}$ with respect to $P(h_{\Sigma}/2)$ lies in the closure of $\Omega$ and the orthogonal projection $p$ of $p_{\max}$ on $P(h_{\Sigma}-h^{\ast})$ lies in the closure of $\Omega(h_{\Sigma}-h^{\ast})$. Up to horizontal translations, we can assume that $p_{\max}$ lies in the axis $\{(0,0)\}\times\mathbb{R}$. We shall show that there exists at least a point $q$ in the closure of $\Omega(h_{\Sigma}-h^{\ast})$ whose distance from $p$ is greater or equal than $r_{d}$. Suppose that this does not occur and denote by $\Sigma^{\ast}$ the connected component of $\Sigma\cap\{h_{\Sigma}-h^{\ast}\leq z\leq h_{\Sigma}\}$ containing $p_{\max}$. Then all the points of $\partial\Sigma^{\ast}=\Sigma^{\ast}\cap P(h_{\Sigma}-h^{\ast})$ are at a distance less than $r_{d}$ from the axis $\{(0,0)\}\times\mathbb{R}$. Consider the round sphere of radius $d/\alpha$ centered at $(0,0,h^{\ast}-d/\alpha)$ in $\mathcal{W}_{f}$, and define $S^{+}$ as its connected component in the upper closed half-space $\{z\geq 0\}$. Then $S^{+}$ is a spherical cap of height $h^{\ast}$ whose boundary is the circle in $P$, centered at the origin with radius $r_{d}$. Let $S^{+}(t)$ be the vertical translation of $S^{+}$ by the vector $te_{3}$. For each $t>h_{\Sigma}$, $S^{+}(t)$ does not touch $\Sigma^{\ast}$. Then, we decrease $t$ and from our assumption, we have that for any $t\in[h_{\Sigma}-h^{\ast},h_{\Sigma}]$ the boundary of $S^{+}(t)$ does not meet the boundary of $\Sigma^{\ast}$. Moreover, for all $t>h_{\Sigma}-h^{\ast}$ there is no interior contact point between $S^{+}(t)$ and $\Sigma^{\ast}$, otherwise, by the maximum principle, $\Sigma$ would coincide with $S^{+}(t)$ and hence it would have a point at height larger than $h_{\Sigma}$. On the other hand, since $p_{\max}$ has height $h_{\Sigma}$, there should be a first interior contact point between $S^{+}(h_{\Sigma}-h^{\ast})$ and $\Sigma^{\ast}$. By the maximum principle, $S^{+}(h_{\Sigma}-h^{\ast})$ and $\Sigma^{\ast}$ must coincide. This is again a contradiction because the points at the boundary of $S^{+}(h_{\Sigma}-h^{\ast})$ are at a distance $r_{d}$ from the axis $\{(0,0)\}\times\mathbb{R}$ while $\Sigma^{\ast}$ has no such points. ∎
@@
Inequality (5.12) guarantees that $\mathcal{L}$ has two loops (see Fig. 5). Using similar arguments as in the proof of Lemma 4.1, we deduce that the smaller loop of $\mathcal{L}$ bounds a closed domain contained in $\Omega(h_{\Sigma}-h^{\ast})$.
@@
Now we denote by $\widetilde{G}(t)$ the image of $\widetilde{G}$ by the translation through the vector $-te_{3}$. There is no values of $t\in[0,h^{\ast}]$ for which the boundary of $\widetilde{G}(t)$ meets $\Sigma$, because the orthogonal projection of the smaller boundary of $\widetilde{G}(t)$ over $P$ is contained in $D(c_{0},r_{0})\subset\Omega$, and the larger boundary of $\widetilde{G}(t)$ is contained in $D(c^{\ast},R^{\ast})\times[h^{\ast},h_{\Sigma}-h^{\ast}]$, which in turn is contained in $\mathcal{O}$. Thus, a possible first contact point between $\widetilde{G}(t_{1})$ and $\Sigma$ for some $t_{1}\in[0,h^{\ast}]$ would be an interior point for both surfaces. If this happens, one has a contradiction by the maximum principle. Let us prove it. Since $G$ is a special catenoid given by $\kappa_{2}=m_{0}k_{1}$, it is inward oriented and since assumption (5.1) implies that $f(t)\geq(1-m_{0})\alpha+m_{0}t$, for all $t\in[\alpha/d,+\infty)$, the curvature diagram of $G$ lies strictly below the curvature diagram corresponding to the set $\mathcal{W}_{f}$ (see Fig. 6, left). Then we can apply Theorem 1.2 at a first interior contact point of $\widetilde{G}(t_{1})$ and $\Sigma$ to conclude that they must coincide, which is a contradiction. Therefore, there is no interior contact point for all $t\in[0,h^{\ast}]$ and $\widetilde{G}(h^{\ast})$ is contained in $\mathcal{O}$. Notice that, the smaller boundary of $\widetilde{G}(h^{\ast})$ is contained in $\Omega$. Moreover, since $r_{0}=r(\Gamma)\leq 1/\Lambda$, then, for any point of $w\in\Gamma$, there is a horizontal translation of $\widetilde{G}(h^{\ast})$ such that its smaller boundary is tangent $\Gamma$ at $w$. Notice that, because of inequality (5.13), by performing such translations, the upper boundary of the translations of $\widetilde{G}(h^{\ast})$ stays inside $\mathcal{O}$. Finally, there cannot be any interior contact point between the horizontal translations of $\widetilde{G}(h^{\ast})$ and $\Sigma$ by the maximum principle. Then the intersection between $\Sigma$ and the region given by $\Omega\times[0,h^{\ast}]$ is empty. ∎
@@
    The Alexandrov reflection method for vertical planes implies that $\Sigma_{1}:=\Sigma\cap\{\frac{h_{\Sigma}}{2}\leq z\leq h_{\Sigma}\}$ is a vertical graph.
@@
and $r_{1}=r_{1}(m_{0},\Gamma,h^{\ast})$ is the radius of the horizontal circle $C_{1}:=G\cap\{z=h^{\ast}\}$, which can be computed through equation (2.1). Moreover, we need a further condition on $g$ that allows to compare $\Sigma\in{\mathcal{W}}_{g}$ with the SW-catenoid $G$ by the maximum principle. Such condition is:
@@
Acknowledgements. The authors would like to thank José Antonio Galvez for stimulating our interest in exploring Ros-Rosenberg theorem in the context of Weingarten surfaces.

### v2

@@ 5 Two Ros-Rosenberg type theorems for SW-surfaces
In the former case, we make a natural assumption on the function $g$ in order to get Ros-Rosenberg result for every surfaces obtained from surfaces in $\mathcal{W}_{g}$ after suitable homotheties, whose ratios can be computed explicitly. We point out that condition (5.1) implies that, when we do a homothety of a surface in $\mathcal{W}_{g},$ its curvature diagram does not intersect a line with slope $m_{0}$ passing through the origin. This allows us to compare our surface with a suitable catenoid by the tangency principle (see Section 1.3).
@@
In view of Remark 5.3, we may suppose that $\Gamma$ is not a circle, that is, $\Lambda>\lambda>0$.
@@
Let us fix a generic unit horizontal vector $v$ and let $Z$ be the horizontal cylinder in $\mathcal{W}_{f}$ whose revolution axis is the straight line orthogonal to $v$, contained in the plane $\{z=d/\beta\}$ and passing through $(c_{\Gamma},d/\beta)$. In this way, the plane $P$ is tangent to $Z$, and the maximum height of $Z$ with respect to $P$ is equals to $2d/\beta$. If $Q_{v}$ is the vertical plane orthogonal to $v$ passing through $(c_{\Gamma},d/\beta)$, we define $\widetilde{Z}$ as the part of $Z$ lying below $Q_{v}$ with respect to $v$, and $\widetilde{Z}(t):=\widetilde{Z}-tv$ as the image of $\widetilde{Z}$ by the translation with respect to the vector $-tv$.
For values $t\gg 0$, $\widetilde{Z}(t)$ does not touch $\Sigma$. By decreasing $t$, $\widetilde{Z}(t)$ approaches $\Sigma$ and by the tangency principle, there cannot be a first contact point between $\widetilde{Z}(t)$ and $\Sigma$ for $t>\omega(\Gamma)$. This implies that $\Sigma$ is contained in the upper (with respect to $v$) closed half-space bounded by $Q_{v}-(\omega(\Gamma)+d/\beta)v$. As $v$ was generically chosen, we deduce that $\Sigma$ is contained in $D((c_{\Gamma},0),\omega(\Gamma)+d/\beta)\times[0,h_{\Sigma}]$.
@@
This yields that $\omega(\Gamma)+d/\beta\leq 1/\lambda+d/\beta<d/\alpha$ and then, the cylinder $\partial S^{+}\times[0,+\infty)$ strictly contains the truncated solid cylinder $D((c_{\Gamma},0),\omega(\Gamma)+d/\beta)\times[0,h_{\Sigma}]$ in its interior. Define $S(\tau):=S^{+}+\tau e_{3}$ to be the translation of $S^{+}$ by the vector $\tau e_{3}$. For any $\tau>2d/\beta,$ the surfaces $S(\tau)$ and $\Sigma$ do not meet. By decreasing $\tau$, if there is a first contact point between $S(\tau_{0})$ and $\Sigma$, for some $\tau_{0}\geq 0$, such a contact point must be necessarily an interior tangency point, since the domain bounded by the orthogonal projection of the boundary of $S(\tau_{0})$ on $P$ strictly contains $\Gamma$. The existence of such an interior contact point is a contradiction by the tangency principle. Therefore, $\Sigma$ must be strictly below $S(0)$.
Since $1/\lambda<d/\alpha$, we can translate $S(0)$ horizontally in such a way that its boundary touch every point of $\Gamma$ we desire from outside. Again by the tangency principle, this implies that $\Sigma$ is contained in the solid cylinder $\overline{\Omega}\times[0,\infty)$. Now, it is easy to prove, using the Alexandrov reflection method with horizontal planes, that $\Sigma$ is a graph over $\Omega$ and therefore a topological disk.
@@
Let $p_{\max}\in\Sigma$ be a point at maximal height and let $\mathcal{O}$ denote the open region bounded by $\Sigma\cup\Omega$. By doing the Alexandrov reflection method with horizontal planes, we get that the reflection of $p_{\max}$ with respect any plane $P(t)$ for $h_{\Sigma}/2<t<h_{\Sigma}$ lies in $\mathcal{O}$. This means the reflection of $p_{\max}$ with respect to $P(h_{\Sigma}/2)$ lies in the closure of $\Omega$ and the orthogonal projection $p$ of $p_{\max}$ on $P(h_{\Sigma}-h^{\ast})$ lies in the closure of $\Omega(h_{\Sigma}-h^{\ast})$. Up to horizontal translations, we can assume that $p_{\max}$ lies in the axis $\{(0,0)\}\times\mathbb{R}$. We shall show that there exists at least a point $q$ in the closure of $\Omega(h_{\Sigma}-h^{\ast})$ whose distance from $p$ is greater or equal than $r_{d}$. Suppose that this does not occur and denote by $\Sigma^{\ast}$ the connected component of $\Sigma\cap\{h_{\Sigma}-h^{\ast}\leq z\leq h_{\Sigma}\}$ containing $p_{\max}$. Then all the points of $\partial\Sigma^{\ast}=\Sigma^{\ast}\cap P(h_{\Sigma}-h^{\ast})$ are at a distance less than $r_{d}$ from the axis $\{(0,0)\}\times\mathbb{R}$. Consider the round sphere of radius $d/\alpha$ centered at $(0,0,h^{\ast}-d/\alpha)$ in $\mathcal{W}_{f}$, and define $S^{+}$ as its connected component in the upper closed half-space $\{z\geq 0\}$. Then $S^{+}$ is a spherical cap of height $h^{\ast}$ whose boundary is the circle in $P$, centered at the origin with radius $r_{d}$. Let $S^{+}(t)$ be the vertical translation of $S^{+}$ by the vector $te_{3}$. For each $t>h_{\Sigma}$, $S^{+}(t)$ does not touch $\Sigma^{\ast}$. Then, we decrease $t$ and from our assumption, we have that for any $t\in[h_{\Sigma}-h^{\ast},h_{\Sigma}]$ the boundary of $S^{+}(t)$ does not meet the boundary of $\Sigma^{\ast}$. Moreover, for all $t>h_{\Sigma}-h^{\ast}$ there is no interior contact point between $S^{+}(t)$ and $\Sigma^{\ast}$, otherwise, by the tangency principle, $\Sigma$ would coincide with $S^{+}(t)$ and hence it would have a point at height larger than $h_{\Sigma}$. On the other hand, since $p_{\max}$ has height $h_{\Sigma}$, there should be a first interior contact point between $S^{+}(h_{\Sigma}-h^{\ast})$ and $\Sigma^{\ast}$. By the tangency principle, $S^{+}(h_{\Sigma}-h^{\ast})$ and $\Sigma^{\ast}$ must coincide. This is again a contradiction because the points at the boundary of $S^{+}(h_{\Sigma}-h^{\ast})$ are at a distance $r_{d}$ from the axis $\{(0,0)\}\times\mathbb{R}$ while $\Sigma^{\ast}$ has no such points. ∎
@@
Inequality (5.12) guarantees that $\mathcal{L}$ has two loops (see Fig. 6). Using similar arguments as in the proof of Lemma 4.1, we deduce that the smaller loop $\mathcal{L}^{\prime}$ of $\mathcal{L}$ bounds a closed domain contained in $\Omega(h_{\Sigma}-h^{\ast})$.
@@
Now we denote by $\widetilde{G}(t)$ the image of $\widetilde{G}$ by the translation through the vector $-te_{3}$. There is no values of $t\in[0,h^{\ast}]$ for which the boundary of $\widetilde{G}(t)$ meets $\Sigma$, because the orthogonal projection of the smaller boundary of $\widetilde{G}(t)$ over $P$ is contained in $D(c_{0},r_{0})\subset\Omega$, and the larger boundary of $\widetilde{G}(t)$ is contained in $D(c^{\ast},R^{\ast})\times[h^{\ast},h_{\Sigma}-h^{\ast}]$, which in turn is contained in $\mathcal{O}$. Thus, a possible first contact point between $\widetilde{G}(t_{1})$ and $\Sigma$ for some $t_{1}\in[0,h^{\ast}]$ would be an interior point for both surfaces. If this happens, one has a contradiction by the tangency principle. Let us prove it. Since $G$ is a special catenoid given by $\kappa_{2}=m_{0}k_{1}$, it is inward oriented and since assumption (5.1) implies that $f(t)\geq(1-m_{0})\alpha+m_{0}t$, for all $t\in[\alpha/d,+\infty)$, the curvature diagram of $G$ lies strictly below the curvature diagram corresponding to the set $\mathcal{W}_{f}$ (see Fig. 7, left). Then we can apply the tangency principle at a first interior contact point of $\widetilde{G}(t_{1})$ and $\Sigma$ to conclude that they must coincide, which is a contradiction. Therefore, there is no interior contact point for all $t\in[0,h^{\ast}]$ and $\widetilde{G}(h^{\ast})$ is contained in $\mathcal{O}$. Notice that, the smaller boundary of $\widetilde{G}(h^{\ast})$ is contained in $\Omega$. Moreover, since $r_{0}=r(\Gamma)\leq 1/\Lambda$, then, for any point of $w\in\Gamma$, there is a horizontal translation of $\widetilde{G}(h^{\ast})$ such that its smaller boundary is tangent $\Gamma$ at $w$. Notice that, because of inequality (5.13), by performing such translations, the upper boundary of the translations of $\widetilde{G}(h^{\ast})$ stays inside $\mathcal{O}$. Finally, there cannot be any interior contact point between the horizontal translations of $\widetilde{G}(h^{\ast})$ and $\Sigma$ by the tangency principle. Then the intersection between $\Sigma$ and the region given by $\Omega\times[0,h^{\ast}]$ is empty. ∎
@@
    The Alexandrov reflection method for horizontal planes implies that $\Sigma_{1}:=\Sigma\cap\{\frac{h_{\Sigma}}{2}\leq z\leq h_{\Sigma}\}$ is a vertical graph.
@@
and $r_{1}=r_{1}(m_{0},\Gamma,h^{\ast})$ is the radius of the horizontal circle $C_{1}:=G\cap\{z=h^{\ast}\}$, which can be computed through equation (2.1). Moreover, we need a further condition on $g$ that allows to compare $\Sigma\in{\mathcal{W}}_{g}$ with the SW-catenoid $G$ by the tangency principle. Such condition is:
@@
Acknowledgements. The authors would like to thank José Antonio Gálvez for stimulating our interest in exploring Ros-Rosenberg theorem in the context of Weingarten surfaces, Isabel Fernández for insightful suggestions and the anonymous referee for reading carefully the article and its valuable remarks. The authors were partially supported by PRIN-2022AP8HZ9, PRIN-20225J97H5 and INdAM-GNSAGA.